# Lessons from pendulums: A design comparison of three lab activities

### v1

@@ VI Reflections and Discussion
As we discussed in the introduction, the purpose of our work here is to wade into the gap between our theoretical perspectives and our curriculum design, in order to make sense of how and why we arrive at different designs despite what we consider to be a shared starting point. Like Boudreaux & Elby [boudreaux_how_2020], the inspiration for our comparison, we compare our designs with the hope of elucidating how our theoretical orientations shape the curricula. While they sought to connect the differences in their animating theoretical frameworks to the differences in their designs, we have sought to make sense of the differences in our design despite the broad alignment of our theoretical perspectives.
Before we move on to make sense of these design differences, we should also take a moment to point out the many similarities between them, many of which reflect our theoretical alignment. On a general level, we all aspire to promote student epistemic agency within the process of experimentation and to center the substance of student ideas. In this we have taken significant efforts to make sure the experiments in our lab are not exercises in confirming or verifying concepts from lecture. Additionally, each of our pendulum activities takes up three class meetings. Furthermore, an important aspect of each of our designs is to disrupt problematic student framing, especially “doing school”. We all expect instructors to be responsive to the substance of student ideas. We each build into our labs the generation and use data representations, and we highlight the need for some way to quantify uncertainty.
In many respects, these shared features of our designs reflect a shared understanding of how framing can influence the agentive element of student behavior, of how, when given the chance, student can be capable of productive scientific work, and how a wide variety of contextual cues can tip students into different framings. These similarities and the underlying theoretical connective tissue between them punctuates the motivation for our comparison. Given this context, we now turn to discuss how and why we arrive at designs with such substantive differences.
@@ VI.1 Presentation of Activity
Above we highlight how each of our labs present the activity to students: At UWB, the experimental question is neutral; at Cornell, the experiment is presented as a model-testing activity; at Tufts, students are asked to investigate Galileo’s claims about pendulums. In the section V.1, we connected these differences to variations in the expectations we seek to establish. Even with these variations, we share overarching aim to communicate to students that their ideas and their data are of central importance.
The neutral phrasing of the UWB instructions clearly focuses the experimental work on the students. We expect the references to the simple harmonic model (Cornell) and Galileo (Tufts) to elicit potentially problematic expectations about verifying already-known, authoritative theoretical claims, so that we can confront and disrupt these expectations. Moreover, the reference to these models sets students up to experience a disagreement between theory and experiment, which in turn generates a genuine (for the students) need to engage in experimental practices; that is, working to reduce uncertainties, use and interpret statistical tools to distinguish data sets, and extending experiments to test emergent questions.
@@ VI.2 Structure of Guidance
The instructions at Cornell also have the effect of scaffolding students collecting data that will indicate a theoretical-empirical discrepancy. Tufts takes a different, minimalist approach to written instructions. In sec. V.2, we highlight how our shared understandings of framing, agency, and affect inform our differing decisions about instructions. To an extent, the difference in our approaches to in-lab instructions indicate a difference in theoretical perspective - or perhaps difference in how we understand the instructional implications of theory.
@@
For Hammer, students’ framing the activity as “doing school” is a primary concern that drives many of the decisions affecting the structure of in-lab guidance. For Scherr and Holmes, while this framing certainly presents a concern, they take the approach of using students’ expectations to follow instructions as a way to nudge students towards productive arenas of experimental work and scientific reasoning. That is, students themselves may not initially see the point of making a histogram and so do so because the instructions tell them to, but thinking about how to best make a histogram and comparing and interpreting that histogram is a rich context for reasoning about and with representations.
@@
There is an additional, affective dimension to Scherr’s and Holmes’s decisions around in-lab guidance: the detailed instructions seek to ensure that students do not feel lost or overwhelmed. To be clear, the instructions at Cornell and UWB are more detailed relative to Tufts; they are much less detailed than a “cookbook” lab. Indeed much of their design work here involves writing open-ended prompts that provide guidance to students while also shifting decision-making onto students. There is an “agentive tradeoff” in the level of detail provided.
@@
Additionally, for the minimal instructions at Tufts we might also think of an “affective tradeoff”. That is, the experience of uncertainty that comes with having to figure it all out gives them an opportunity to manage and regulate potential feelings of discomfort - meta-affective learning [radoff_its_2019] is itself a desired outcome. While the in-lab instructions at Cornell and UWB seek to provide some amount of affective support for students, at Tufts much of that responsibility is shifted onto the instructor. In this way, the graduate and undergraduate TAs at Tufts are empowered to be more flexible and adaptable in how they respond to specific students and emergent idiosyncrasies.
Of course, the lab instructors at Cornell and UWB are also empowered and able to respond to students. That said, an aspect of the reasoning for more granular instructions at Cornell and UWB is that they minimize the potential negative influence of inexperienced TAs or instructors with discordant pedagogical beliefs. On the other hand, part of Hammer’s management of the labs at Tufts includes an active emphasis on TA education that aligns with the role of TAs in the curriculum.
@@
Finally, we note a key difference with respect to curricular objectives with respect to statistical analysis. At UWB and Cornell, much of the guidance within the materials concerns the definition and use of the t’. At Tufts, there is no “content” objective in this lab with respect to canonical statistics. Rather, the objective is entirely epistemological, that students recognize and take on the challenge of quantifying uncertainty.
@@ VI.3 Models of Students
Moving forward, one area in particular that this exercise in comparison has provided inspiration for each of us is the degree to which many of our design decisions were informed by our often unstated “working model” of the student population. At Cornell and Tufts, experience has taught Holmes and Hammer that students in their labs will often know of the simple harmonic period equation, while that is not Scherr’s experience at UWB. These considerations were consequential in how we operationalized our theoretical commitments. Scherr and Holmes both have a sense of the affective consequences for their students of a lack of in-lab instructions: in part, the guiding nature of their in-lab instructions are designed to meet students where they are.
We refer to these expectations as a working model to reflect both their unfinished and partially tacit nature, and to highlight these expectations as connected to our theoretical commitments. Implicit in our desire to shift students’ framings is an expectation of how students will think and act coming into our labs. Certainly research has shaped these expectations [hu_examining_2018, hu_examining_2018-1, hu_qualitative_2017, Hu_Zwickl_2017, smith_how_2020, smith_lets_2020, smith_surprise_2018, sundstrom_problematizing_2020, descamps_dynamics_2024, scherr_student_2009], along with our experiences in the classroom. Still, as we engaged in our comparisons, we often found ourselves needing to explicate and clarify our understandings of the students. In reflecting on the ways in which the Cornell and UWB labs structure student engagement with statistical and representational tools, the Tufts instructional team realized we had unstated expectations about students’ capability to generate and use novel representational forms. The lab activity at Tufts was shaped by these expectations, even though they were tacit throughout our design process.
Additionally, in discussing our various expectations of students’ – as individuals – we also recognized how structuring good collaborative groups in our labs remains an open-ended question for each of us. Collaboration is built into our curriculum on a fundamental level, and at each institution we engage in a handful of strategies to promote the emergence of equitable and supportive small groups. Still, there is an element of uncertainty with regards to our expectations about how to best support the development good collaborative groups and what precisely “good” looks like.
@@ VII Closing thoughts
The differences among our lab activities demonstrate that epistemic agency and framing can engender multiple, distinct forms of pedagogical structures. We see this as an opportunity: to innovate, to explore, to try things out, and to conduct scholarship that deepens our understanding of what works. The varied instructional environments studied here certainly do not span the entire design space. Given the heterogeneity of national and global lab contexts, exploration and innovation are necessary for the pursuit of developing flexible and adaptable pedagogical structures or, more broadly, generating actionable design inspiration.
@@
In reflecting on how curriculum developers and the broader PER ecosystem can work towards more useful research outputs, Elby & Yerdelen-Damar introduced the notion of instructionally generative fodder: descriptions of curriculum, related research of student activity and learning in designed contexts, and more general research into learning that provide actionable insight into learning and how to support it [elby_rethinking_2020]. To optimize the generativity of descriptions of curriculum, they suggest that developers make available and clear to instructors the rationale for design decisions, including revisions.
We support that suggestion. Our examination of our own curricula helped us recognize and articulate aspects of the reasoning that play significant roles in our developing and refining activities and materials. With this in mind, and given our work here, we suggest a shift in how developers describe and explain curricula, to focus on design reasoning. This shift puts the emphasis on how decisions are made and how a curriculum is expected to mediate learning. Instructors’ access to developers’ reasoning for particular features will help them in their own reasoning as they adopt and perhaps customize curricula for their students.
@@
Consider how the experimental question is posed in our three pendulum lab activities. Focusing on design decisions, one might say “our understanding of student agency and framing guided our decision about the phrasing of the experimental question.” This explains the decision, but provides no insight into how we arrived at different phrasings. Focusing on the design reasoning requires us to discuss expectations of students (at Cornell and Tufts) to be aware of the pendulum period equation, to describe how these questions lead into an epistemological productive failure and elicit-confront-resolve approach, and to articulate how each of these phrasings work to emphasize the substance of students’ data and experimental work.
@@
A shift to design reasoning echoes the argument that Brookes et al. [brookes_implementing_2020] outline in their paper on the motivations for ISLE. Building from educational theorists MacMillan & Garrison [Garrison_MacMillan], Brookes et al. understand teaching as an “intentional activity in which we strive to change students’ attitudes and beliefs” and go on to “suggest that if we are to make better connections between education theory and classroom implementation, both researchers and implementers need to articulate their underlying intentionalities more explicitly.”
@@
In recent years there has been a push for PER scholars to include more detailed and more useful descriptions of theory in curriculum development. Our comparison further underscores necessity and complexity of this task. We contend that a focus on design reasoning can help make descriptions of curricula more generative for practitioners. Design reasoning can also be generative for researchers: expectations about how a curriculum should mediate learning can be studied! Given that research-based pedagogies are taken up in noticeably different ways (if they are taken up at all) [turpen_construction_2010, dancy_how_2016, affriyenni_navigating], it would be useful for research to highlight the consequential bits of curricula and crucial interactions and dynamics.
@@ Acknowledgements.
We would like to thank the faculty and staff at Cornell, UWB, and Tufts involved in developing and implementing these labs. We acknowledge the Cornell University College of Arts and Sciences Active Learning Initiative for providing support for part of the development of the Thinking Critically in Physics curriculum. We also thanks members of the Tufts STEM Education program, in particular Ira Caspari and students in the Qualitative Research in STEM Education course for feedback on initial comparative analyses.

### v2

@@ VI Reflections and Discussion
As we discussed in the introduction, the purpose of our work here is to wade into the gap between our theoretical perspectives and our curriculum designs. Like Boudreaux & Elby [7], the inspiration for our comparison, we compare our designs with the hope of elucidating how our theoretical orientations shape the curricula. While they sought to connect the differences in their theoretical frameworks to the differences in their designs, we have sought to make sense of the differences in our design despite the alignment of our theoretical perspectives.
First, we review the similarities, many of which reflect our theoretical alignment. We all aspire to promote student epistemic agency within the process of experimentation and to center the substance of student ideas. We all work to make sure the experiments in our labs are not exercises in confirming or verifying concepts from lecture; try to shift students away from problematic framing; expect instructors to be responsive to the substance of student ideas; build into our labs the generation and use of data representations; highlight the need for some way to quantify uncertainty, and, in general, aim at communicating to students that their ideas and their data are of central importance.
These similarities reflect a shared understanding of how framing can influence the agentive element of student behavior, of how, when given the chance, student can be capable of productive scientific work, and how a wide variety of contextual cues can tip students into different framings. These similarities and the underlying theoretical connective tissue between them punctuates the motivation to consider how and why we arrive at designs with such substantive differences.
@@ VI.1 Presentation of Activity
Above, we highlight how each of our labs present the activity to students: At UWB, the experimental question is neutral; at Cornell, the experiment is presented as a model-testing activity; at Tufts, students are asked to investigate the claims of a renowned authority. In the section V.1, we connected these differences to variations in the expectations we seek to cultivate in students.
The neutral phrasing of the UWB instructions clearly focuses the experimental work on the students. We expect the references to the simple harmonic model (Cornell) and Galileo (Tufts) to elicit potentially problematic expectations about verifying already-known, authoritative theoretical claims, so that we can confront and disrupt these expectations. Moreover, the reference to these models sets students up to experience a disagreement between theory and experiment, which in turn generates a genuine (for the students) need to engage in experimental practices of working to reduce uncertainties, use and interpret statistical tools to distinguish data sets, and extend experiments to test emergent questions.
@@ VI.2 Structure of Guidance
The instructions at Cornell also have the effect of scaffolding students’ collecting data that will indicate a theoretical-empirical discrepancy. Tufts takes a different, minimalist approach to written instructions. In sec. V.2, we highlight how our shared understandings of framing, agency, and affect inform our differing decisions about instructions. To an extent, the difference in our approaches to in-lab instructions indicate a difference in theoretical perspective—or perhaps difference in how we understand the instructional implications of theory.
@@
For Hammer, students’ framing the activity as “doing school” [46] is a primary concern that drives many of the decisions affecting the structure of in-lab guidance. For Scherr and Holmes, while this framing certainly presents a concern, they take the approach of using students’ expectations to follow instructions as a way to nudge students towards productive arenas of experimental work and scientific reasoning. That is, students themselves may not initially see the point of making a histogram and so do so because the instructions tell them to, but thinking about how to best make a histogram and comparing and interpreting that histogram is a rich context for reasoning about and with representations.
@@
There is an additional, affective dimension to Scherr’s and Holmes’s decisions around in-lab guidance: the detailed instructions seek to ensure that students do not feel lost or overwhelmed. To be clear, while the instructions at Cornell and UWB are more detailed relative to Tufts, they are much less detailed than a “cookbook” lab. Indeed much of their design work here involves writing open-ended prompts that provide guidance to students while also shifting decision-making onto students.
@@
Across the curricula, the designs reflect different choices over an “agentive tradeoff” in the level of detail to provide. There is also an “affective tradeoff,” with respect to students’ experiences of uncertainty, greater at Tufts with less written guidance. Students learning to manage and regulate potential feelings of discomfort—meta-affective learning [62] is itself a desired outcome. While the in-lab instructions at Cornell and UWB provide support for students with respect to agency and affect, at Tufts that falls much more to the instructor, a graduate or undergraduate TA. In this way, TAs at Tufts are responsible and empowered to be more flexible and adaptable in how they respond to specific students and emergent idiosyncrasies.
Of course, the lab instructors at Cornell and UWB are also empowered and able to respond to students, but the onus is reduced. Part of the reasoning for more granular instructions at Cornell and UWB is that they minimize the potential negative influence of inexperienced TAs or instructors with discordant pedagogical beliefs. Indeed, the Tufts lab design places a heavy demand on TAs, and experience has shown the importance of supporting their learning as instructors. This is another form of tradeoff: Maintaining that support in the face of varying enrollments (and thus numbers of TAs), levels of engagement and skill of both the TAs and the faculty assigned to that role, has proved to be challenging. There has been noticeable variation in the effectiveness of the labs across different semesters.
@@
Finally, we note a key difference with respect to curricular objectives with respect to statistical analysis. At UWB and Cornell, much of the guidance within the materials concerns the definition and use of the t’ statistic. At Tufts, there is no “content” objective in this lab with respect to canonical statistics. Rather, the objective is entirely epistemological, that students recognize and take on the challenge of quantifying uncertainty. The difference in course structure (stand-alone lab course versus component of a lecture course) plays a role in this decision-making: the stand-alone structure affords opportunities to engage in invention activities about statistics.
@@ VI.3 Models of Students
Moving forward, one area in particular where this exercise in comparison has provided inspiration for each of us is the degree to which our design decisions are informed by often unstated “working models” of the students. At Cornell and Tufts, experience has taught Holmes and Hammer that students in their labs will have heard of both Galileo and the simple harmonic period equation, while that is not Scherr’s experience at UWB. These considerations were consequential in how we operationalized our theoretical commitments. Scherr and Holmes both have a sense of the affective consequences for their students of a lack of in-lab instructions: in part, the guiding nature of their in-lab instructions is designed to meet students where they are.
We refer to these expectations as working models to reflect both their unfinished and partially tacit nature, and to highlight these expectations as connected to our theoretical commitments. Implicit in our desire to shift students’ framings is an expectation of how students will think and act coming into our labs. Research has shaped our expectations [42, 41, 40, 43, 79, 76, 77, 82, 14, 71], along with our experiences in and out of the classroom. Still, as we engaged in our comparisons, we often found ourselves needing to explicate and clarify our understandings of the students. In reflecting on the ways in which the Cornell and UWB labs structure student engagement with statistical and representational tools, the Tufts instructional team realized the lab design was shaped by unstated expectations about students’ abilities to generate and use novel representational forms. This motivated a redesign of the third lab session to include more explicit attention to data representation and analysis.
Additionally, in discussing our various expectations of students – as individuals – we also recognized how structuring collaborative groups remains an open question. Collaboration is built into our curriculum on a fundamental level, and at each school, we engage in a handful of strategies to promote the emergence of equitable and supportive small groups. Still, we are all uncertain how to best to support groups’ working effectively, as well as how to assess "effectiveness" of collaboration.
@@ VII Closing thoughts
The differences among our lab activities demonstrate that a theoretical framework of epistemic agency and framing can engender multiple, distinct forms of pedagogical structures.Our labs provide a sample, but they certainly do not span the entire design space. There is plenty of opportunity, for curriculum designers as for instructors, to innovate, to explore, to try things out, as well as to conduct scholarship that deepens our understanding of what works. Given the heterogeneity of national and global lab contexts [36, 24], there is not only opportunity but also necessity, for local inquiry by instructors, as well as for the development of curricula that are flexible and adaptable to local circumstances.
@@
In reflecting on how curriculum developers and the broader PER ecosystem can work towards more useful research outputs, Elby & Yerdelen-Damar introduced the notion of instructionally generative fodder: descriptions of curriculum, related research of student activity and learning in designed contexts, and more general research that provides actionable insight into learning and how to support it [18]. To optimize the generativity of curricular products, they suggest that developers make the rationale for design decisions available and clear to instructors, including points of uncertainty.
In recent years, there has been a push for PER scholars to include more detailed and more useful descriptions of theory in curriculum development. In doing this work, researchers have commonly focused on design decisions and choices—making more decisions explicit and connecting them to theoretical frameworks. We suggest a slight shift in focus, to instead emphasize design reasoning. That is, descriptions of curriculum should bring to the fore the reasoning involved in how design decisions are made and how a curriculum is expected to mediate learning. This suggestion is largely a rephrasing and synthesis of Elby & Yerdelen-Damar’s [18] discussion of generativity in curricular products; it also echoes other arguments that it is important for teachers to understand the purposes and aims of curricular materials [74, 75, 5, 13, 30, 65].
@@
Consider how the experimental question is posed in our three pendulum lab activities. Focusing on design decisions, one might say that our understanding of student agency and framing guided our decisions about how to phrase the experimental question. This explains the decision, but provides no insight into how we arrived at different phrasings. Focusing on the design reasoning requires us to discuss expectations of students (at Cornell and Tufts) to be aware of the pendulum period equation, to describe how these questions lead into an epistemologically productive failure and elicit-confront-resolve approach, and to articulate how each of these phrasings work to emphasize the substance of students’ data and experimental work.
@@
A shift to design reasoning also echoes the argument that Brookes et al. [9] outline in their paper on the motivations for ISLE. Building from educational theorists MacMillan & Garrison [51], they “suggest that if we are to make better connections between education theory and classroom implementation, both researchers and implementers need to articulate their underlying intentionalities more explicitly.”
Robertson et al. [65] illustrate the utility of knowledge of what we are calling design reasoning in a study of Learning Assistants in a university-level physics course. They specifically provide evidence that “knowledge of the purposes of questions or sequences of questions within [research-based instructional materials] […] can support (or serve as a resource in) the enactment of responsive instruction” (p. 151). Robertson et al. build from a long tradition of education scholars who maintain that making visible the "pedagogical judgments" [13] and "hidden decisions" [30] of curricular materials can support teachers in enacting theoretically oriented and research-based curricula [74, 75, 5, 64]. We extend this claim to recommend that curricular materials and research on curriculum development in PER ought to foreground design reasoning.
@@
Our examination of our own curricula helped us recognize and articulate aspects of the reasoning that play significant roles in our developing and refining activities and materials. We contend that a focus on design reasoning can help make descriptions of curricula more generative for practitioners. Design reasoning can also be generative for researchers: expectations about how a curriculum will support learning should inform studies of what takes place. Given that research-based pedagogies are taken up in noticeably different ways (if they are taken up at all) [87, 12, 1], it would be useful for research to highlight the consequential aspects of curricula and crucial interactions.
@@ Acknowledgements.
We would like to thank the faculty and staff at Cornell, UWB, and Tufts involved in developing and implementing these labs. We acknowledge the Cornell University College of Arts and Sciences Active Learning Initiative for providing support for part of the development of the Thinking Critically in Physics curriculum. We also thanks members of the Tufts STEM Education program, in particular Sophia Jeon, Miguel Vasquez-Vega, Ira Caspari, and the students in the Qualitative Research in STEM Education course for feedback on initial comparative analyses. Finally, we appreciate the thorough feedback of our anonymous reviewers, in particular the connections to Robertson et al. [65]